OpenStudy (anonymous):

y=x^3+x^2-9x-9, list the intercepts and test for symmetry

OpenStudy (campbell_st):

use the factor theorem P(-1) = 0 so (x + 1) is a factor then by polynomial division the factors are (x + 1)(x^2 - 9) or (x +1)(x-3)(x+3) then the x-intercepts are x = -3, -1, 3 the y intercept is when x = 0 or y = -9 test to see if the function is odd, even of neither P(-x) = -x^3 + x^2 + 9x - 9 the function is neither odd nor even so does not have symmetry